OpenStudy (anonymous):

(4/5)^-4 please explain the steps please

OpenStudy (amistre64):

what steps?

OpenStudy (amistre64):

its impossible to explain steps to something that has no instructions attached to it. what is it your spose to do to this thing?

OpenStudy (anonymous):

solving it?

OpenStudy (amistre64):

theres nothing to solve ..

OpenStudy (amistre64):

its like asking: 3 show the steps and solve it ...

OpenStudy (anonymous):

when ever you have a negative power flip the denominator and numerator. then your power becomes positive. (4/5)^(-5) = (5/4) ^ (5) = 5^5/4^5 = 3125/1024 when you have a power for the whole fraction, raise the denominator and numerator to the same power

OpenStudy (anonymous):

thank u so much that makes alot of sense

OpenStudy (anonymous):

you are most welcome. BTW attractive profile picture. :D

OpenStudy (anonymous):

ty